OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (2, 1) m = 0

OpenStudy (ivettef365):

you have slope of 0, so that is a line horizontally that goes on y=1

OpenStudy (anonymous):

y2-y1=m(x2-x1)+b, 1-y=0(2-1)+b, 1-y=b, -y=b-1, y=-b+1, y=1-b I think thats right